OpenStudy (anonymous):

Solve the system by using a table. -0.2x-0.3y=0.2 0.4x-0.3y=1.4

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

do you know how to make a table?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

isn't it (2,2)?

jimthompson5910 (jim_thompson5910):

let me check

jimthompson5910 (jim_thompson5910):

-0.2x-0.3y=0.2 0.4x-0.3y=1.4 multiply the top equation by -1, then add the equations to get 0.2x+0.3y=-0.2 0.4x-0.3y=1.4 --------------- 0.6x+0y = 1.2 so that means, 0.6x = 1.2 x = ??

OpenStudy (anonymous):

2

OpenStudy (anonymous):

so, i'm right?

jimthompson5910 (jim_thompson5910):

x = 2, is correct now use this to find y

jimthompson5910 (jim_thompson5910):

y = 2 is NOT correct

OpenStudy (anonymous):

1?

jimthompson5910 (jim_thompson5910):

0.4x-0.3y=1.4 0.4*2-0.3y=1.4 0.8 - 0.3y = 1.4 -0.3y = 1.4 - 0.8 -0.3y = 0.6 y = ???